# Graph Transformers Dream of Electric Flow

### v1

@@ 6 Learning Positional Encoding for Molecular Regression
To test this, we evaluate the performance of our proposed linear Transformer on a molecular regression task on two real-world datasets: QM9 (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014) and ZINC (Irwin et al., 2012). The regression target is constrained solubility (Dwivedi et al., 2023). Our experiments are based on the Graph Transformer (GT) implementation from Dwivedi and Bresson (2020). In Table 2, we compare three loss values: GT without PE, GT with Laplacian Eigenvector as PE (LapPE), and GT with Linear Transformer output as PE. We choose to use LapPE as a baseline as it was the original PE of choice in Dwivedi and Bresson (2020). We present details of the datasets in Appendix 12.1. The architecture of linear Transformer is a modified version of (5.1), detailed in Appendix 12.2. Other experiment details, including precise definitions of each model, can be found in Appendix 12.3.
@@
The difference between GT and GT + LapPE is substantial for both QM9 and ZINC, highlighting the importance of PE in both cases.
Going from GT + LapPE to GT + Linear Transformer for ZINC, there is a significant further improvement in loss (a drop of 0.063). This is remarkable, considering that the linear Transformer accounts for less than $0.7\%$ of the total number of parameters. Note that the SOTA error for ZINC regression is significantly lower than 0.138 on more recent architectures (see e.g. Lim et al. (2022)); the significance of our result is in demonstrating the improvement just by replacing the LapPE with the linear Transformer, while keeping everything else fixed.
In contrast, when going from GT + LapPE to GT + Linear Transformer for QM9, the difference is essentially zero. We conjecture that this may be because QM9 molecules (average of about $9$ nodes and $19$ edges) are considerably simpler than ZINC (average of about $23$ nodes and $40$ edges). Thus there may not be too many additional useful features to learn, and LapPE close to optimal. It is still consistent with our theory, for the Linear Transformer to do as well as LapPE.
7 Acknowledgement
The authors would like to thank Professor Xavier Bresson from the National University of Singapore for providing the GitHub code which facilitated some of the experiments in this work. Suvrit Sra acknowledges generous support from the Alexander von Humboldt Foundation.

### v2

@@ 6 Learning Positional Encoding for Molecular Regression
To test this, we evaluate the performance of our proposed linear Transformer on a molecular regression task on two real-world datasets: QM9 (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014) and ZINC (Irwin et al., 2012). The regression target is constrained solubility (Dwivedi et al., 2023). Our experiments are based on the Graph Transformer (GT) implementation from Dwivedi & Bresson (2020). In Table 2, we compare three loss values: GT without PE, GT with Laplacian Eigenvector as PE (LapPE), and GT with Linear Transformer output as PE. We use LapPE as a baseline because it was the PE used in Dwivedi & Bresson (2020). We present details of the datasets in Appendix 14.1. The linear Transformer architecture is a modified version of (7), detailed in Appendix 14.2. Other experiment details, including precise model definitions, can be found in Appendix 14.3.
@@
The difference between GT and GT + LapPE is substantial for both QM9 and ZINC, highlighting the importance of PE in both cases. Going from GT + LapPE to GT + Linear Transformer for ZINC, there is a significant further improvement in loss (about $30\%$). This is remarkable, considering that the linear Transformer accounts for less than $0.7\%$ of the total number of parameters. Note that the SOTA error for ZINC regression is significantly lower than 0.138 on more recent architectures; the significance of our result is in demonstrating the improvement just by replacing the LapPE with the linear Transformer, while keeping everything else fixed. In contrast, going from GT + LapPE to GT + Linear Transformer for QM9, the difference is essentially zero. We conjecture that this may be because QM9 molecules (average of about $9$ nodes and $19$ edges) are considerably smaller than ZINC (average of about $23$ nodes and $40$ edges). Thus there may not be too many additional useful features to learn, and LapPE close to optimal. It is still consistent with our theory, for the Linear Transformer to do as well as LapPE.
7 Ethics Statement
The authors have adhered to ICLR’s code of ethics during the writing of this paper, and when conducting the research described herein. There are no ethics concerns which need to be highlighted.
8 Reproducibility
We use both synthetic data and open-source molecular datasets in our experiments. Details of our experiment procedures and implementations have been provided in Section 3.5, Section 4.1, Section 6 and Appendix 14.
9 Acknowledgements
Suvrit Sra Acknowledges generous support from the Alexander von Humboldt Foundation’s AI Professorship. The authors thank Aaron Schild for the many valuable discussions, and Professor Xavier Bresson from the National University of Singapore for sharing the code which facilitated some of the experiments in this work.